# KramaBench: A Benchmark for AI Systems on Data Intensive Tasks

### v2

@@ 4.1 Ablation Studies
| Rows Sampled | 10 | 50 | 100 | 150 |
| Overall Performance (%) | 22.89 | 24.68 | 23.36 | 22.58 |
| Tokens (Mean) | 14,077.2 | 37,592.3 | 64,548.9 | 92,116.1 |
@@
Retrieval Mechanisms. Using Oracle input for DS-Guru improves the performance for the overall dataset across all domains and LLMs (by 9.98% on average and up to 20%) except for GPT-o3, Claude-3.5, and DeepSeek-R1 on wildfire (Table 5). These results under the design of DS-Guru show that supplying samples of the gold files can lead to more successful pipelines. We also studied the sensitivity of OPS against the sample size from each file. Table 6 shows that the performance of the system does not meaningfully increase with larger samples.
@@
The benefits of the Oracle in smolagents DR shows the same trend (improvements of around 10%), suggesting that agentic retrieval is not qualitatively closer to perfect retrieval than OPS in terms of file extraction. Even with the Oracle input, the agentic smolagents DR with out-of-the-box LLMs still struggle to solve a lot of the tasks (59% overall with Claude-3.7). These results point to weaknesses in data-dependent reasoning (e.g., pipeline design), in addition to extracting the right files.
| System | Models | Combined | Legal | Wildfire | ||||||
| Full | Oracle | Obscured | Full | Oracle | Obscured | Full | Oracle | Obscured | ||
| DS-Guru no-context | GPT-o3 | 12.54% | 10.72% | 11.15% | 5.08% | 6.45% | 9.46% | 20.00% | 15.00% | 13.07% |
| GPT-4o | 7.19% | 2.52% | 8.60% | 4.37% | 5.04% | 9.83% | 10.00% | 0.002% | 7.21% | |
| Claude-3-5 | 8.50% | 4.85% | 9.93% | 6.99% | 4.64% | 11.30% | 10.00% | 0.00% | 8.37% | |
| DS-Guru one-shot | GPT-o3 | 12.73% | 26.98% | 8.99% | 15.47% | 24.89% | 11.62% | 10.00% | 29.08% | 5.99% |
| GPT-4o | 26.03% | 27.45% | 11.15% | 12.06% | 14.89% | 10.27% | 40.00% | 40.00% | 12.16% | |
| Claude-3-5 | 19.56% | 18.82% | 6.11% | 5.03% | 8.56% | 7.35% | 34.08% | 29.07% | 4.70% | |
| DS-Guru few-shot | GPT-o3 | 7.08% | 44.74% | 20.40% | 4.18% | 40.41% | 20.29% | 10.00% | 49.08% | 20.52% |
| GPT-4o | 21.92% | 34.21% | 11.90% | 8.85% | 28.41% | 15.47% | 35.00% | 40.00% | 7.82% | |
| Claude-3-5 | 25.06% | 23.02% | 13.46% | 6.04% | 16.98% | 12.18% | 44.08% | 29.07% | 14.92% | |
| Smolagents DR | GPT-o3 | 40% | 45% | 20% | 50% | 30% | 30% | 30% | 60% | 10% |
| GPT-4o | 50% | 40% | 30% | 50% | 40% | 30% | 50% | 40% | 30% | |
| Claude-3-5 | 55% | 60% | 30% | 40% | 50% | 20% | 70% | 70% | 40% | |
@@
Data Leakage. To study to what extent different systems are solving tasks via external knowledge present in previous knowledge data instead of producing a reliable data pipeline, we manually curated obscured inputs for 20% of tasks in KramaBench, where some data fields are changed such that a correct pipeline would still produce a correct solution, but a system relying on memorization cannot. For example, in a query spanning multiple locations, the real place names may be swapped for fictional ones, i.e., Los Angeles might be changed to “La-La Land."
@@
For both smolagents DR and DS-Guru few-shot, the performance under the obscured input is 12-16% lower compared to the oracle input (Table 7). Interestingly, compared to Full input, Obscured input improved the performance for DS-Guru but significantly degraded smolagent for the legal workload. These observations and the stark difference between the Full and Obscured input performances on wildfire suggest two distinctive plausible explanations for our observations: (1) Prior knowledge could discourage attempts at data-dependent reasoning. (2) Prior knowledge could be serving as an unintended reward signal in agentic data-dependent reasoning, which possibly can either improve or reduce the performance of the system. These effects are important topics for future research.
| Models | |||||||
| Variant | Automation setting | GPT-o3 | GPT-4o | Claude-3.5 | Llama3-3Instruct | DeepSeek-R1 | Qwen2-5Coder |
| DS-GURU no-context | End-to-end automation | 9.64% | 1.62% | 7.45% | 1.19% | 3.14% | 3.72% |
| Pipeline Design | 40.60% | 30.83% | 31.06% | 26.74% | 18.94% | 27.35% | |
| Sub-task Implementation | 12.95% | 9.27% | 10.65% | 8.28% | 12.08% | 7.52% | |
| DS-GURU one-shot | End-to-end automation | 20.80% | 7.61% | 10.85% | 4.81% | 6.35% | 6.43% |
| Pipeline Design | 42.14% | 19.75% | 25.49% | 19.24% | 10.60% | 22.19% | |
| Sub-task Implementation | 17.24% | 11.42% | 10.12% | 7.83% | 11.37% | 10.38% | |
| DS-GURU few-shot | End-to-end automation | 22.08% | 8.28% | 14.35% | 4.48% | 6.35% | 9.98% |
| Pipeline Design | 41.58% | 16.67% | 29.46% | 16.83% | 6.44% | 14.65% | |
| Sub-task Implementation | 19.75% | 13.67% | 16.14% | 8.87% | 10.89% | 12.09% | |
@@
Diversity of Abilities Required from LLM Agents. Tested independently, both pipeline design (+19.50% GPT-o3) and sub-task implementation (+5.39% GPT-4o) substantially outperformed end-to-end automation. In addition, we observe that LLMs also have varying capability profiles: GPT-o3 is strong at high-level pipeline design (42%) but weak at implementing those pipelines (20%); interestingly, it scores higher on end-to-end automation (22%) than on some implementation tasks. DeepSeek-R1 exhibits the opposite pattern (6.5% on pipeline design vs 11% on implementation). These patterns provide strong evidence that single-agent approaches are insufficiently reliable for real-world data science, as success depends on multiple heterogeneous skills, such as robust parsing of noisy inputs, query parsing and planning, identifying and performing data-cleaning/ transformation, coding, and iterative debugging.
@@ 4.2 Deeper Dive: Failure Analysis
The challenge with beach identification is that the "Beach Name" column encodes both the beach and sampling location. Cliff Pond (DCR) @ Main refers to the Main (street) sampling location of the Cliff Pond beach (Figure 2). Facing many near-duplicate files in the data lake, systems do not have a clear global schema or geographical domain knowledge that they could use to understand this encoding scheme. As a result, both DS-Guru and smolagents DR failed on this sub-task, despite smolagents DR’s agentic control flow. This case highlights the need to incorporate prior knowledge and discover clarifications about under-specified conventions from the data (Mao et al., 2019).
Towards this end, we analyzed the traces of DS-Guru (few-shot, GPT-o3 & Claude 3.5) with the agentic system diagnosis framework proposed in Cemri et al. (2025). Respectively 24% (GPT-o3) and 43% (Claude 3.5) of all 104 tasks suffer from "failure to ask for clarification" and thus were not solved correctly. However, KramaBench expects that a human expert could solve the tasks without additional clarifications by exploring and understanding the data. In this example, a human could decipher the beach name conventions with common U.S. geographical knowledge. Reasoning models currently lack similar capabilities to gain holistic understandings of the input data and fail to incorporate prior knowledge.
@@ 5 Related Work
Evaluations of LLM-Powered Agentic Systems. Benchmarks for question answering (QA) have shifted toward evaluating agentic solutions. These benchmarks require iterative retrieval, query parsing, planning, tool use, and temporal awareness. Recent works include FanOutQA (Zhu et al., 2024), MultiHop-RAG (Tang & Yang, 2024), CRAG (Yang et al., 2024), BrowseComp Wei et al. (2025), which test end-to-end retrieval systems, MEQA (Li et al., 2024) for multi-hop reasoning with explanation chains, and MINTQA (He et al., 2024) for scaffolding long knowledge. These tasks differ from data science tasks, as they only require information retrieval and joins, but no data-intensive processing. Benchmarks such as DS-1000 (Lai et al., 2023), DA-Code (Huang et al., 2024), ARCADE (Yin et al., 2023), DataSciBench (Zhang et al., 2025a), DSEval (Zhang et al., 2024c) focus instead on implementing detailed instructions in general programming languages, specifically in data science tasks, differentiating themselves from other benchmarks like SWE-Bench (Jimenez et al., 2024), ML-Bench (Tang et al., 2024b), BigCodeBench (Zhuo et al., 2025). More recently, new benchmarks such as DSBench (Jing et al., 2025) and BLADE (Gu et al., 2024) have started to evaluate the ability to create an implementation plan. Benchmarks like ScienceAgentBench (Chen et al., 2025) and BixBench (Mitchener et al., 2025) evaluate using domain knowledge. Although such benchmarks assess specific capabilities, they fall short of capturing the full complexity of real-world data science pipelines.
@@ 6 Conclusion
We provide full artifacts—including code, data, workloads, and evaluation scripts—via our public repository at https://github.com/mitdbg/Kramabench. The main paper section 2 and Appendix C describe the process obtained to design and curate the task based on the datasets for each domain. Scripts to reproduce these steps can be found in the main repository. All datasets, benchmark frameworks, benchmark curation semi-automation scripts, reference pipelines and other accompanying annotations, and our reference system DS-Guru are available in our repository. The experimental analysis of different system under test in Section 4 can be reproduced using Python scripts also available in the public repository.
@@ Appendix A Extended Experiment Results
| Archaeology | Astronomy | Biomedical | Environment | Legal | Wildfire | Overall | ||
| DS-Guru no-context | GPT-o3 | 25% | 1.73% | 3.50% | 1.35% | 3.35% | 24.87% | 9.64% |
| GPT-4o | 0.00% | 1.41% | 1.98% | 0.45% | 1.46% | 1.45% | 1.62% | |
| Claude-3.5 | 16.67% | 1.62% | 2.87% | 1.17% | 7.33% | 13.63% | 7.45% | |
| Llama3-3Instruct | 0.00% | 1.43% | 1.70% | 0.98% | 1.37% | 1.44% | 1.19% | |
| DeepSeek-R1 | 0.00% | 1.50% | 2.49% | 2.60% | 1.61% | 6.46% | 3.14% | |
| Qwen2-5Coder | 0.00% | 1.37% | 2.02% | 1.07% | 1.44% | 13.68% | 3.72% | |
| DS-Guru one-shot | GPT-o3 | 25% | 3.00% | 8.63% | 7.66% | 19.15% | 45.95% | 20.80% |
| GPT-4o | 8.33% | 1.40% | 9.38% | 2.60% | 2.74% | 19.39% | 7.61% | |
| Claude-3.5 | 0.00% | 4.15% | 2.15% | 6.21% | 6.68% | 34.99% | 10.85% | |
| Llama3-3Instruct | 0.00% | 1.42% | 10.38% | 0.98% | 5.48% | 9.81% | 4.81% | |
| DeepSeek-R1 | 0.00% | 1.57% | 3.39% | 2.60% | 8.30% | 14.81% | 6.35% | |
| Qwen2-5Coder | 0.00% | 1.36% | 2.22% | 12.59% | 1.15% | 16.48% | 6.43% | |
| DS-Guru few-shot | GPT-o3 | 25% | 3.53% | 8.95% | 19.6% | 13.89% | 50.73% | 22.08% |
| GPT-4o | 16.67% | 2.76% | 8.97% | 2.60% | 2.80% | 17.18% | 8.28% | |
| Claude-3.5 | 16.67% | 1.52% | 1.96% | 11.21% | 7.01% | 39.16% | 14.35% | |
| Llama3-3Instruct | 0.00% | 1.35% | 6.98% | 0.93% | 2.15% | 14.49% | 4.48% | |
| DeepSeek-R1 | 8.33% | 2.64% | 2.87% | 19.08% | 8.39% | 30.29% | 6.34% | |
| Qwen2-5Coder | 8.33% | 2.40% | 4.35% | 12.64% | 9.06% | 16.48% | 9.98% | |
| smolagents DR | GPT-o3 | 41.67% | 16.67% | 33.33% | 50% | 50% | 38.1% | 41.36% |
| GPT-4o | 33.33% | 0.00% | 11.11% | 35% | 40% | 38.1% | 30.77% | |
| Claude-3-5 | 33.33% | 0.00% | 22.22% | 60% | 46.67% | 52.38% | 41.35% | |
| Claude-3-7 | 33.33% | 16.67% | 44.44% | 60% | 63.33% | 52.38% | 50% | |

### v3

@@ 4.1 Ablation Studies
Retrieval Mechanisms. Overall, using Oracle input for DS-Guru improves the performance across all domains and LLMs (by on average and up to ) except for GPT-o3 for which in the no-context scenario no significant improvement is registered (Table 5). These results under the design of DS-Guru show that supplying samples of the gold files can lead to more successful pipelines. We also studied the sensitivity of OPS against the sample size from each file. Table 7 shows that the performance of the system does not meaningfully increase with larger samples.
@@
The benefits of the Oracle in smolagents DR show the same trend, suggesting that agentic retrieval is not qualitatively closer to perfect retrieval than OPS in terms of file extraction. Even with the Oracle input, the agentic smolagents DR with out-of-the-box LLMs still struggle to solve a large portion of the tasks ( overall with Claude-3.7). These results point to weaknesses in data-dependent reasoning (e.g., pipeline design), in addition to extracting the right files.
| Rows | Overall | Tokens |
| Sampled | Score () | (Mean) |
| 10 | 22.89 | 14,077.2 |
| 50 | 24.68 | 37,592.3 |
| 100 | 23.36 | 64,548.9 |
| 150 | 22.58 | 92,116.1 |
| System | Model | Full | Oracle | Obscured |
| DS-Guru no-context | GPT-o3 | 5.87±0.71 | 5.36±1.20 | 3.69±1.63 |
| Claude-3.7 | 1.88±0.81 | 4.11±0.79 | 0.00±0.00 | |
| DS-Guru one-shot | GPT-o3 | 16.67±2.93 | 21.35±0.84 | 7.85±1.27 |
| Claude-3.7 | 1.31±1.22 | 3.27±1.31 | 0.50±0.50 | |
| DS-Guru few-shot | GPT-o3 | 24.98±1.25 | 43.71±1.94 | 23.02±0.49 |
| Claude-3.7 | 8.29±1.12 | 19.75±3.36 | 3.80±3.80 | |
| smolagents DR | GPT-o3 | 28.07±8.80 | 25.86±4.57 | 10.07±0.30 |
| Claude-3.7 | 55.83±3.41 | 60.67±1.23 | 12.77±0.00 | |
| smolagents Reflexion | GPT-o3 | 30.53±10.79 | 32.33±4.44 | 9.52±3.95 |
| Claude-3.7 | 55.37 ± 3.36 | 62.81±3.10 | 13.94±0.00 | |
| smolagents PDT | GPT-o3 | 7.57±1.28 | 7.69±2.69 | 3.87±3.52 |
| Claude-3.7 | 12.01±1.03 | 14.38±2.15 | 2.23±0.00 |
@@
Data Leakage. To study to what extent different systems are solving tasks via external knowledge present in previous knowledge data instead of producing a reliable data pipeline, we manually curated obscured inputs for all tasks in KramaBench, where data fields are changed such that a correct pipeline would still produce a correct solution, but a system relying on memorization cannot. For example, in a query spanning multiple locations, the real place names may be swapped for fictional ones, i.e., Los Angeles might be changed to “La-La Land."
@@
For both smolagents single DR and DS-Guru few-shot, the performance under the obscured input is lower compared to the regular and oracle inputs (Table 7). Interestingly, Claude as a backend model seems the most sensitive to the obscured workloads, e.g. with a sharp drop from to in the Reflexion baseline, suggesting that the solution of the benchmark in these variants may heavily rely on parametric knowledge. These observations and the stark difference between the Full and Obscured input performances suggest two distinctive plausible explanations for our observations: (1) Prior knowledge could discourage attempts at data-dependent reasoning. (2) Prior knowledge could be serving as an unintended reward signal in agentic data-dependent reasoning, which possibly can either reduce the performance of the system.
Cross-domain Accuracy Difference The per-domain accuracy of the best performing system (smolagents Reflexion) in KramaBench varies as much as on Archeology and on Wildfire. Our analysis of system traces show that the primary source of these cross-domain accuracy differences is the differences in the types of data task challenges that each domain emphasize on. We discuss this in more detail in Subsection D.1.
| Models | |||||
| Variant | Automation setting | GPT-o3 | GPT-4o | Llama3 Instruct | DeepSeek-R1 |
| DS-Guru no-context | End-to-end automation | 5.87 ± 0.71 | 0.00 ± 0.00 | 0.33 ± 0.46 | 0.33 ± 0.46 |
| Pipeline Design | 39.75 ± 0.73 | 29.73 ± 1.16 | 21.80 ± 1.30 | 0.63 ± 0.47 | |
| Sub-task Implementation | 10.69 ± 0.35 | 3.32 ± 0.54 | 3.88 ± 0.25 | 1.48 ± 0.08 | |
| DS-Guru one-shot | End-to-end automation | 16.67 ± 2.93 | 6.08 ± 1.25 | 3.42 ± 0.37 | 0.00 ± 0.00 |
| Pipeline Design | 41.71 ± 1.00 | 21.26 ± 0.64 | 15.53 ± 0.73 | 1.59 ± 0.77 | |
| Sub-task Implementation | 17.33 ± 0.48 | 5.08 ± 0.57 | 2.80 ± 0.08 | 3.91 ± 0.44 | |
| DS-Guru few-shot | End-to-end automation | 24.98 ± 1.25 | 8.67 ± 1.48 | 8.40 ± 1.01 | 0.33 ± 0.46 |
| Pipeline Design | 37.99 ± 2.64 | 20.31 ± 0.07 | 12.81 ± 1.04 | 1.11 ± 0.72 | |
| Sub-task Implementation | 22.05 ± 0.71 | 6.60 ± 0.29 | 5.53 ± 0.49 | 5.79 ± 0.72 | |
@@
Diversity of Abilities Required from LLM Agents. Tested independently, pipeline design and subtask implementation outperformed end-to-end automation, e.g, even in the no-context scenario, using GPT-o3 as a model, the pipeline design score is compared to end-to-end, In addition, we observe that LLMs also have varying capability profiles: GPT-o3 is strong at high-level pipeline design () but weak at implementing those pipelines (); interestingly, it scores higher on end-to-end automation () than on some implementation tasks. Although with worse absolute results, DeepSeek-R1 exhibits the opposite pattern ( on pipeline design vs on implementation). These patterns provide strong evidence that single-model approaches are insufficiently reliable for real-world data science, as success depends on multiple heterogeneous skills, such as robust parsing of noisy inputs, query parsing and planning, identifying and performing data-cleaning/ transformation, coding, and iterative debugging.
@@ 4.2 Deeper Dive: Failure Analysis
The challenge with beach identification is that the "Beach Name" column encodes both the beach and sampling location. Cliff Pond (DCR) @ Main refers to the Main (street) sampling location of the Cliff Pond beach (Figure 5). Facing many near-duplicate files in the data lake, systems do not have a clear global schema or geographical domain knowledge that they could use to understand this encoding scheme. As a result, both DS-Guru and smolagents DR failed on this sub-task, despite smolagents DR’s agentic control flow. This case highlights the need to incorporate prior knowledge and discover clarifications about under-specified conventions from the data (Mao et al., 2019). Towards this end, we analyzed the traces of DS-Guru (few-shot, GPT-o3 & Claude 3.5) with the agentic system diagnosis framework proposed in Cemri et al. (2025). Respectively (GPT-o3) and (Claude 3.5) of all 104 tasks suffer from "failure to ask for clarification" and thus were not solved correctly. However, KramaBench expects that a human expert could solve the tasks without additional clarifications by exploring and understanding the data. In this example, a human could decipher the beach name conventions with common U.S. geographical knowledge. Reasoning models currently lack similar capabilities to gain holistic understandings of the input data and fail to incorporate prior knowledge.
@@ 5 Related Work
Evaluations of LLM-Powered Agentic Systems. Benchmarks for question answering (QA) have shifted toward evaluating agentic solutions. These benchmarks require iterative retrieval, query parsing, planning, tool use, and temporal awareness. Recent works include FanOutQA (Zhu et al., 2024), MultiHop-RAG (Tang and Yang, 2024), CRAG (Yang et al., 2024), BrowseComp (Wei et al., 2025), which test end-to-end retrieval systems, MEQA (Li et al., 2024) for multi-hop reasoning with explanation chains, and MINTQA (He et al., 2024) for scaffolding long knowledge. These tasks differ from data science tasks, as they only require information retrieval and joins, but no data-intensive processing. Benchmarks such as DS-1000 (Lai et al., 2023), DA-Code (Huang et al., 2024), ARCADE (Yin et al., 2023), DataSciBench (Zhang et al., 2025a), DSEval (Zhang et al., 2024c) focus instead on implementing detailed instructions in general programming languages, specifically in data science tasks, differentiating themselves from other benchmarks like SWE-Bench (Jimenez et al., 2024), ML-Bench (Tang et al., 2024b), BigCodeBench (Zhuo et al., 2025). More recently, new benchmarks such as DSBench (Jing et al., 2025) and BLADE (Gu et al., 2024) have started to evaluate the ability to create an implementation plan. Benchmarks like ScienceAgentBench (Chen et al., 2025) and BixBench (Mitchener et al., 2025) evaluate using domain knowledge. Although such benchmarks assess specific capabilities, they fall short of capturing the full complexity of real-world data science pipelines.
@@ 6 Conclusion
We provide full artifacts—including code, data, workloads, and evaluation scripts—via our public repository at https://github.com/mitdbg/Kramabench. The main paper section 2 and Appendix D describe the process obtained to design and curate the task based on the datasets for each domain. Scripts to reproduce these steps can be found in the main repository. All datasets, benchmark frameworks, benchmark curation semi-automation scripts, reference pipelines and other accompanying annotations, and our reference system DS-Guru are available in our repository. The experimental analysis of different system under test in Section 4 can be reproduced using Python scripts also available in the public repository.
Acknowledgments
We are grateful for the support from the DARPA ASKEM Award HR00112220042, the ARPA-H Biomedical Data Fabric project, NSF DBI 2327954, a grant from Liberty Mutual, a Google Research Award, and the Amazon Research Award. Additionally, our work has been supported by contributions from Amazon, Google, and Intel as part of the MIT Data Systems and AI Lab (DSAIL) at MIT, along with NSF IIS 1900933. This research was sponsored by the United States Air Force Research Laboratory and the Department of the Air Force Artificial Intelligence Accelerator and was accomplished under Cooperative Agreement Number FA8750-19-2-1000. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the Department of the Air Force or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein.
@@ Appendix A Extended Experiment Results
| Archeology | Astronomy | Biomedical | Environment | Legal | Wildfire | Overall | ||
| DS-Guru no-context | GPT-o3 | 16.67 ± 00.00 | 0.00 ± 00.00 | 0.10 ± 0.14 | 5.00 ± 4.08 | 0.00 ± 00.00 | 14.18 ± 4.28 | 5.87 ± 0.71 |
| GPT-4o | 0.00 ± 00.00 | 0.00 ± 00.00 | 0.00 ± 00.00 | 0.00 ± 00.00 | 0.00 ± 00.00 | 0.00 ± 00.00 | 0.00 ± 00.00 | |
| Llama-3.3-Instruct | 0.00 ± 00.00 | 0.00 ± 00.00 | 0.00 ± 00.00 | 0.00 ± 00.00 | 0.00 ± 00.00 | 1.59 ± 2.24 | 0.33 ± 0.46 | |
| Deepseek-R1 | 2.78 ± 3.93 | 0.00 ± 00.00 | 0.00 ± 00.00 | 0.00 ± 00.00 | 0.00 ± 00.00 | 0.00 ± 00.00 | 0.33 ± 0.46 | |
| Qwen2-5Coder* | 0.00 | 1.37 | 2.02 | 1.07 | 1.44 | 13.68 | 3.72 | |
| Claude-3.5* | 16.67 | 1.62 | 2.87 | 1.17 | 7.33 | 13.63 | 7.45 | |
| Claude-3.7 | 2.78 ± 3.93 | 3.33 ± 4.71 | 0.00 ± 00.00 | 0.00 ± 00.00 | 1.11 ± 1.57 | 4.38 ± 3.92 | 1.88 ± 0.81 | |
| DS-Guru one-shot | GPT-o3 | 19.44 ± 3.93 | 0.00 ± 00.00 | 0.53 ± 0.75 | 16.11 ± 9.26 | 10.00 ± 00.00 | 40.01 ± 4.22 | 16.67 ± 2.93 |
| GPT-4o | 8.33 ± 6.80 | 6.67 ± 4.71 | 3.70 ± 00.00 | 0.00 ± 00.00 | 4.41 ± 2.17 | 13.69 ± 3.06 | 6.08 ± 1.25 | |
| Llama-3.3-Instruct | 16.67 ± 00.00 | 0.00 ± 00.00 | 0.10 ± 0.14 | 0.00 ± 00.00 | 0.74 ± 0.52 | 5.97 ± 2.55 | 3.42 ± 0.37 | |
| Deepseek-R1 | 0.00 ± 00.00 | 0.00 ± 00.00 | 0.00 ± 00.00 | 0.00 ± 00.00 | 0.00 ± 00.00 | 0.00 ± 00.00 | 0.00 ± 00.00 | |
| Qwen2-5 Coder* | 0.00 | 1.36 | 2.22 | 12.59 | 1.15 | 16.48 | 6.43 | |
| Claude-3.5* | 0.00 | 4.15 | 2.15 | 6.21 | 6.68 | 34.99 | 10.85 | |
| Claude-3.7 | 2.78 ± 3.93 | 0.00 ± 00.00 | 0.00 ± 00.00 | 3.33 ± 2.36 | 0.00 ± 00.00 | 1.59 ± 2.24 | 1.31 ± 1.22 | |
| DS-Guru few-shot | GPT-o3 | 13.89 ± 3.93 | 0.00 ± 00.00 | 0.10 ± 0.14 | 49.56 ± 0.87 | 9.26 ± 2.92 | 52.92 ± 0.72 | 24.98 ± 1.25 |
| GPT-4o | 13.89 ± 3.93 | 3.33 ± 4.71 | 0.00 ± 00.00 | 0.00 ± 00.00 | 4.84 ± 3.16 | 25.69 ± 1.15 | 8.67 ± 1.48 | |
| Llama-3.3-Instruct | 22.22 ± 3.93 | 0.00 ± 00.00 | 0.20 ± 0.14 | 0.00 ± 00.00 | 3.33 ± 00.00 | 23.25 ± 3.06 | 8.40 ± 1.01 | |
| Deepseek-R1 | 0.00 ± 00.00 | 0.00 ± 00.00 | 0.00 ± 00.00 | 0.00 ± 00.00 | 0.00 ± 00.00 | 1.59 ± 2.24 | 0.33 ± 0.46 | |
| Qwen2-5Coder* | 8.33 | 2.40 | 4.35 | 12.64 | 9.06 | 16.48 | 9.98 | |
| Claude-3.5* | 16.67 | 1.52 | 1.96 | 11.21 | 7.01 | 39.16 | 14.35 | |
| Claude-3.7 | 5.56 ± 3.93 | 0.00 ± 00.00 | 0.00 ± 00.00 | 15.00 ± 4.08 | 8.89 ± 6.29 | 10.09 ± 9.67 | 8.29 ± 1.12 | |
| smolagents DR | GPT-o3 | 8.33 ± 6.80 | 0.00 ± 00.00 | 0.00 ± 00.00 | 1.42 ± 1.09 | 4.51 ± 0.84 | 22.01 ± 1.39 | 7.57 ± 1.28 |
| GPT-4o* | 33.33 | 0.00 | 11.11 | 35.00 | 40.00 | 38.10 | 30.77 | |
| Claude-3.5* | 33.33 | 0.00 | 22.22 | 60.00 | 46.67 | 52.38 | 41.35 | |
| Claude-3.7 | 44.44 ± 3.93 | 50.00 ± 8.16 | 38.89 ± 7.86 | 60.56 ± 7.74 | 61.23 ± 12.70 | 60.16 ± 7.43 | 55.83 ± 3.41 | |
| smolagents Reflexion | GPT-o3 | 30.56 ± 10.39 | 46.67 ± 4.71 | 29.63 ± 5.24 | 23.89 ± 14.55 | 26.67 ± 14.40 | 35.04 ± 8.36 | 30.53 ± 10.79 |
| Claude-3.7 | 41.67 ± 0.00 | 45.00 ± 5.00 | 50.00 ± 27.78 | 56.25 ± 8.75 | 58.33 ± 1.67 | 65.38 ± 0.86 | 55.37 ± 3.36 | |
| smolagents PDT | GPT-o3 | 8.33 ± 6.80 | 0.00 ± 00.00 | 0.00 ± 00.00 | 1.42 ± 1.09 | 4.51 ± 0.84 | 22.01 ± 1.39 | 7.57 ± 1.28 |
| Claude-3.7 | 19.44 ± 7.86 | 6.67 ± 4.71 | 5.56 ± 5.56 | 9.84 ± 2.86 | 6.25 ± 4.95 | 22.58 ± 4.05 | 12.01 ± 1.03 | |